Slove the system of equatoins by addition method
-4x=3y-10
-8x-10y= -28

Answer by checkley77(12844)

-4x=3y-10 multiply by -2 & add.
-8x-10y=-28
8x+6y=20
--------------------
-4y=-8
y=-8/-4
y=2 ans.
-4x=3*2-10
-4x=6-10
-4x=-4
x=-4/-4
x=1 ans.
Proof:
-8*1-10*2=-28
-8-20=-28
-28=-28